# Squeezed Quasinormal Modes from Nonlinear Gravitational Effects

### v1

@@ III Degree of Squeezing
that suggests that squeezing of the sub-Poissonian nature Mandel (1982). In Appendix B, we also map out a perturbatively exact analysis based on Refs . Ou (1994); Li and Kumar (1994); Suhara et al. (1996). For coherent state initial conditions for the fundamental mode and second harmonic mode in vacuum, where $\langle\hat{X}^{2}_{j\omega}(0)\rangle=\langle\hat{P}^{2}_{j\omega}(0)\rangle=\frac{1}{2}$, one obtains that,
$\langle\hat{X}^{2}_{j\omega}(\zeta)\rangle=\langle\hat{X}^{2}_{j\omega}(0)\rangle S_{j\omega X},~~\text{and}~~\langle\hat{P}^{2}_{j\omega}(\zeta)\rangle=\langle\hat{P}^{2}_{j\omega}(0)\rangle S_{j\omega P},~~j=1,2.$ (36)
Above, $\zeta=\frac{1}{\sqrt{2}}\lambda|\langle A_{\omega}(0)\rangle|t=\frac{1}{\sqrt{2}}\lambda|\alpha_{\omega}(0)|t$ and,
$S_{\omega X}$ $=$ $(1-\zeta\tanh\zeta)^{2}\text{sech}^{2}\zeta+2\tanh^{2}\zeta\text{sech}^{2}\zeta,$
$S_{2\omega X}$ $=$ $\frac{1}{2}(\tanh\zeta+\zeta\text{sech}^{2}\zeta)^{2}+\text{sech}^{4}\zeta,$
$S_{\omega P}$ $=$ $\frac{1}{2}(\sinh\zeta+\zeta\text{sech}\zeta)^{2}+\text{sech}^{2}\zeta,$
$S_{2\omega P}$ $=$ $2\tanh^{2}\zeta+(1-\zeta\tanh\zeta)^{2},$ (37)
which reduces to Eq. (35) at small $\zeta$. For $\zeta\gg 1$, one obtains that, $S_{\omega X}\rightarrow 4\zeta^{2}e^{-2\zeta}$, $S_{2\omega X}\rightarrow 1/2$, $S_{\omega P}\rightarrow e^{2\zeta}/8$, and $S_{2\omega P}\rightarrow\zeta^{2}$ Ou (1994); Li and Kumar (1994); Suhara et al. (1996). This suggests substantial squeezing can in principle be generated through this mechanism, especially an exponential enhancement in the squeezing of the $\hat{P}$ quadrature of the fundamental mode.
$|\zeta|=\frac{1}{\sqrt{2}}\lambda|\alpha_{\omega}(0)|t\approx\sqrt{\frac{16\pi G\hbar\omega^{2}}{c^{5}}}\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right||\alpha_{\omega}(0)|,$ (38)
$|\zeta|=\sqrt{\frac{16\pi G\hbar\omega^{2}}{c^{5}}}\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right|\sqrt{\frac{c^{5}}{32\pi G\hbar\omega^{2}}}\approx\frac{1}{\sqrt{2}}\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right|h_{\omega}\approx 0.1h_{\omega},$ (39)
since $\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right|\sim 0.15$. While the amplitude of the gravitational radiation is substantially small at our detectors, $h_{\omega}^{D}\approx 10^{-22}$, at the source it can be substantially high, and can be of the order of unity, suggesting that $|\zeta|\approx 0.1$. Assuming that the fundamental mode is in a coherent state, we are also in the regime of validity of Eq. (35),
$\langle(\Delta\hat{X}_{\omega})^{2}\rangle$ $\approx$ $\frac{1}{2}-\lambda^{2}t^{2}|\alpha_{\omega}|^{2}/4=\frac{1}{2}\left(1-|\zeta|^{2}\right)=0.495,~~\text{and}$
$\langle(\Delta\hat{P}_{\omega})^{2}\rangle$ $\approx$ $\frac{1}{2}+\lambda^{2}t^{2}|\alpha_{\omega}|^{2}/4=\frac{1}{2}\left(1+|\zeta|^{2}\right)=0.505.$ (40)
For reasons of concreteness and simplicity, we focused on a scenario where the fundamental mode begins in a coherent state. The merger dynamics may well prepare them in an acoherent quantum state, however. We defer a detailed analysis of this to future work. Also, we initialized the second harmonic mode in its vacuum state, but the merger dynamics could also excite the second harmonic mode. The resulting dynamics of the fundamental mode can then be understood in terms of a parametric oscillator, a natural scenario for describing quantum mechanical squeezing Chaturvedi et al. (2002); Mallet et al. (2011); Murch et al. (2013); Wu et al. (1987); Adamyan et al. (2015); Manikandan and Qvarfort (2023). We discuss the mapping to a parametrically excited oscillator for the fundamental mode in an Appendix. C.
IV Discussion
We have estimated the degree of squeezing generated in long-lived quasinormal modes from gravitational self-interactions. The estimates are valid for the perturbative regime following the merger event. For this scenario, where only weak nonlinearities are involved, we find that higher harmonic generation of quasinormal modes leads to $\sim 1\%$ squeezing in the fundamental mode. As mentioned immediately above, our simplifying assumptions about initial conditions are not unlikely to be overly pessimistic. More generally, earlier stages of radiation close to merger, where nonlinearities are stronger, can be expected to yield larger effects. Recent work suggests possibilities for testing these ideas experimentally Manikandan and Wilczek (2025a, , b); Parikh et al. (2020, 2021b, 2021a); Tobar et al. (2024); Shenderov et al. (2024).
V Acknowledgments
FW is supported by the U.S. Department of Energy under grant Contract Number DE-SC0012567 and by the Swedish Research Council under Contract No. 335-2014-7424. SKM was supported in part by the Swedish Research Council under Contract No. 335-2014-7424 and in part by the Wallenberg Initiative on Networks and Quantum Information (WINQ). We thank Maulik Parikh and Igor Pikovski for stimulating discussions around these subjects.
Or alternatively, the number of gravitons,
Based on this, we can estimate the amplitude ratio,
The preceding calculations also allow us to estimate the effective coupling $\lambda$ as,
$H_{I}$ $\approx$ $-\sqrt{2}\Lambda\left[\hat{h}_{j,0}(t)\right]^{2}\hat{h}_{j}^{(p)}(t)$ (45)
$=$ $-\hbar\sqrt{\hbar}\Lambda\omega^{3}\left(\frac{16\pi G}{c^{5}}\right)^{3/2}(A_{\omega}e^{-i\omega t}+A_{\omega}^{\dagger}e^{i\omega t})^{2}(A_{2\omega}e^{-i2\omega t}+A_{2\omega}^{\dagger}e^{i2\omega t})$
where,
$\lambda=2\sqrt{\hbar}\Lambda\left(\frac{16\pi G\omega^{6}}{c^{5}}\right)^{3/2}.$ (46)
Appendix B Exact analysis of squeezing in second harmonic generation
We closely follow the analysis in Ref. Li and Kumar (1994); Ou (1994). Consider the dynamical equations that describe second harmonic generation,
Define quantum operators about the classical mean-field solutions,
@@ Appendix B Exact analysis of squeezing in second harmonic generation
One can now define the following change of variables,
$u_{\omega}=\frac{|\langle A_{\omega}(t)\rangle|}{|\langle A_{\omega}(0)\rangle|},~~u_{2\omega}=\frac{\sqrt{2}|\langle A_{2\omega}(t)\rangle|}{|\langle A_{\omega}(0)\rangle|},~~\text{and}~~\zeta=\frac{1}{\sqrt{2}}\lambda|\langle A_{\omega}(0)\rangle|t,$ (51)
@@
$\frac{du_{\omega}}{d\zeta}=-u_{\omega}u_{2\omega},~~\frac{du_{2\omega}}{d\zeta}=u_{\omega}^{2},$ (52)
where perfect phase-matching is assumed, $\text{Arg}(\langle A_{2\omega}(t)\rangle)=\phi_{2\omega}(t)=2\text{Arg}(\langle A_{\omega}(t)\rangle)=2\phi_{\omega}(t).$ Hence, only the real amplitudes come into play. The system of equations admit solutions (assuming initial conditions $u_{\omega}(0)=1$, and $u_{2\omega}(0)=0$),
$u_{\omega}(\zeta)=\text{sech}\zeta,~~u_{2\omega}(\zeta)=\tanh\zeta.$ (53)
For the generalized position and momentum quadratures, $x_{j\omega}(\zeta)=\frac{1}{\sqrt{2}}(a_{j\omega}e^{-i\phi_{j\omega}}+a^{\dagger}_{j\omega}e^{i\phi_{j\omega}})$ and $p_{j\omega}(\zeta)=\frac{1}{i\sqrt{2}}(a_{j\omega}e^{-i\phi_{j\omega}}-a^{\dagger}_{j\omega}e^{i\phi_{j\omega}}),$ one obtains equations,
$\frac{d\hat{X}_{\omega}}{d\zeta}$ $=$ $-\hat{X}_{\omega}\tanh\zeta-\sqrt{2}\hat{X}_{2\omega}\text{sech}\zeta,$
$\frac{d\hat{X}_{2\omega}}{d\zeta}$ $=$ $\sqrt{2}\hat{X}_{\omega}\text{sech}\zeta$
$\frac{d\hat{P}_{\omega}}{d\zeta}$ $=$ $\hat{P}_{\omega}\tanh\zeta-\sqrt{2}\hat{X}_{2\omega}\text{sech}\zeta,$
$\frac{d\hat{P}_{2\omega}}{d\zeta}$ $=$ $\sqrt{2}\hat{P}_{\omega}\text{sech}\zeta.$ (54)
Solutions are straightforward. One obtains,
$\hat{X}_{\omega}(\zeta)$ $=$ $(1-\zeta\tanh\zeta)\hat{X}_{\omega}(0)\text{sech}\zeta-\sqrt{2}\hat{X}_{2\omega}(0)\tanh\zeta\text{sech}\zeta,$
$\hat{X}_{2\omega}(\zeta)$ $=$ $\frac{1}{\sqrt{2}}(\tanh\zeta+\zeta\text{sech}^{2}\zeta)\hat{X}_{\omega}(0)+\hat{X}_{2\omega}(0)\text{sech}^{2}\zeta,$
$\hat{P}_{\omega}(\zeta)$ $=$ $\hat{P}_{\omega}(0)\text{sech}\zeta-\frac{1}{\sqrt{2}}(\sinh\zeta+\zeta\text{sech}\zeta)\hat{P}_{2\omega}(0),$
$\hat{P}_{2\omega}(\zeta)$ $=$ $\sqrt{2}\hat{P}_{\omega}(0)\tanh\zeta+(1-\zeta\tanh\zeta)\hat{P}_{2\omega}(0).$ (55)
@@
$\langle\hat{X}^{2}_{j\omega}(\zeta)\rangle=\langle\hat{X}^{2}_{j\omega}(0)\rangle S_{j\omega X},~~\text{and}~~\langle\hat{P}^{2}_{j\omega}(\zeta)\rangle=\langle\hat{P}^{2}_{j\omega}(0)\rangle S_{j\omega P},$ (56)
where
$S_{\omega X}$ $=$ $(1-\zeta\tanh\zeta)^{2}\text{sech}^{2}\zeta+2\tanh^{2}\zeta\text{sech}^{2}\zeta,$
$S_{2\omega X}$ $=$ $\frac{1}{2}(\tanh\zeta+\zeta\text{sech}^{2}\zeta)^{2}+\text{sech}^{4}\zeta,$
$S_{\omega P}$ $=$ $\frac{1}{2}(\sinh\zeta+\zeta\text{sech}\zeta)^{2}+\text{sech}^{2}\zeta,$
$S_{2\omega P}$ $=$ $2\tanh^{2}\zeta+(1-\zeta\tanh\zeta)^{2}.$ (57)
For large $\zeta$, note that, $S_{\omega X}\rightarrow 4\zeta^{2}e^{-2\zeta}$, $S_{2\omega X}\rightarrow 1/2$, $S_{\omega P}\rightarrow e^{2\zeta}/8$, and $S_{2\omega P}\rightarrow\zeta^{2}$ Ou (1994); Li and Kumar (1994); Suhara et al. (1996), suggesting the exponential enhancement in the noise of the phase quadrature. While the predictions are limited by the validity of the linearized equations that, for instance, assume a highly intense coherent field on the fundamental mode to start with, we can take these results to indicate that substantial squeezing can, in principle, be generated via the second harmonic generation in the fundamental mode. However, for the scenarios we are interested in, we have also excluded the decay, or damping, of the quasinormal modes, so our predictions are relevant for weakly damped quasinormal modes that are long-lived, applicable at timescales much smaller than the lifetime of quasinormal modes. Hence, the relevant degree of squeezing can be estimated based on the limit of small $\zeta$, captured already by the perturbative expansion, Eq. (35) of the main text.
Appendix C Connections to a parametrically excited oscillator
The squeezing of the fundamental mode can also be understood in terms of a parametric oscillator if the second harmonic mode is populated to begin with. To see this, note that the following Lagrangian,
$L_{p}=\frac{1}{2}\left(\dot{h}_{44}^{(p)}\right)^{2}-\frac{\omega_{440}^{2}}{2}\left(h_{440}^{(p)}\right)^{2}+\sqrt{2}\Lambda h_{220}^{2}h_{440}^{(p)},$ (58)
yields our Eq. (22) for the second harmonic field,
$\ddot{h}_{44}^{(p)}+\omega_{440}^{2}h_{440}^{(p)}=\sqrt{2}\Lambda h_{220}^{2}.$ (59)
The Lagrangian for the free fundamental mode is given by,
$L_{0}=\frac{1}{2}\dot{h}_{220}^{2}-\frac{\omega_{220}^{2}}{2}h_{220}^{2},$ (60)
which in the limit $\Lambda=0$ generates the wave equation for the fundamental mode,
$\ddot{h}_{220}+\omega_{220}^{2}h_{220}=0.$ (61)
However, with nonzero $\Lambda$, variation of the total Lagrangian $L=L_{p}+L_{0}$ with respect to $\dot{h}_{220}$ yields,
$\ddot{h}_{220}+\left[\omega_{220}^{2}-2\sqrt{2}\Lambda h_{44}^{(p)}\right]h_{220}=0.$ (62)
We see that the mode $h_{440}^{(p)}$ modulates the frequency of the fundamental mode in a time-dependent manner. Since the temporal part of $h_{44}^{(p)}$ oscillates at frequency $\omega_{p}=2\omega_{220}$, we can approximate it as $h_{44}^{(p)}(t)\approx h_{0}^{(p)}\cos(2\omega_{220}t)$. With this, we note that the equation governing the fundamental mode is modified as,
where $\epsilon=\sqrt{2}\Lambda h_{0}^{(p)}$. This is Mathieu’s equation in the standard form that describes a parametrically driven oscillator with $\omega_{p}=2\omega_{220}$. Such a drive naturally squeezes the fundamental mode as it can be understood as time-dependent modulations of the quantum harmonic potential, which has a squeezing effect Chaturvedi et al. (2002); Mallet et al. (2011); Murch et al. (2013); Wu et al. (1987); Adamyan et al. (2015); Manikandan and Qvarfort (2023).

### v2

@@ Appendix C Exact analysis of squeezing in second harmonic generation
Following the standard practice, one can define the quantum operators about the classical mean-field solutions,
@@
One can now define the following change of variables Li and Kumar (1994); Ou (1994),
$u_{\omega}=\frac{|\langle A_{\omega}(t)\rangle|}{|\langle A_{\omega}(0)\rangle|},~~u_{2\omega}=\frac{\sqrt{2}|\langle A_{2\omega}(t)\rangle|}{|\langle A_{\omega}(0)\rangle|},~~\text{and}~~\tau=\frac{1}{\sqrt{2}}\lambda|\langle A_{\omega}(0)\rangle|t,$ (67)
@@
$\frac{du_{\omega}}{d\tau}=-u_{\omega}u_{2\omega},~~\frac{du_{2\omega}}{d\tau}=u_{\omega}^{2},$ (68)
where perfect phase-matching is assumed, $\text{Arg}(\langle A_{2\omega}(t)\rangle)=\phi_{2\omega}(t)=2\text{Arg}(\langle A_{\omega}(t)\rangle)=2\phi_{\omega}(t).$ Hence, only the real amplitudes come into play. The system of equations admit solutions (assuming initial conditions $u_{\omega}(0)=1$, and $u_{2\omega}(0)=0$Li and Kumar (1994); Ou (1994),
$u_{\omega}(\tau)=\text{sech}\tau,~~u_{2\omega}(\tau)=\tanh\tau.$ (69)
For the generalized position and momentum quadratures, $\hat{X}_{j\omega}(\tau)=\frac{1}{\sqrt{2}}(a_{j\omega}e^{-i\phi_{j\omega}}+a^{\dagger}_{j\omega}e^{i\phi_{j\omega}})$ and $\hat{P}_{j\omega}(\tau)=\frac{1}{i\sqrt{2}}(a_{j\omega}e^{-i\phi_{j\omega}}-a^{\dagger}_{j\omega}e^{i\phi_{j\omega}}),$ one obtains equations Li and Kumar (1994); Ou (1994),
$\frac{d\hat{X}_{\omega}}{d\tau}$ $=$ $-\hat{X}_{\omega}\tanh\tau-\sqrt{2}\hat{X}_{2\omega}\text{sech}\tau,$
$\frac{d\hat{X}_{2\omega}}{d\tau}$ $=$ $\sqrt{2}\hat{X}_{\omega}\text{sech}\tau$
$\frac{d\hat{P}_{\omega}}{d\tau}$ $=$ $\hat{P}_{\omega}\tanh\tau-\sqrt{2}\hat{X}_{2\omega}\text{sech}\tau,$
$\frac{d\hat{P}_{2\omega}}{d\tau}$ $=$ $\sqrt{2}\hat{P}_{\omega}\text{sech}\tau.$ (70)
Solutions are straightforward. One obtains Li and Kumar (1994); Ou (1994),
$\hat{X}_{\omega}(\tau)$ $=$ $(1-\tau\tanh\tau)\hat{X}_{\omega}(0)\text{sech}\tau-\sqrt{2}\hat{X}_{2\omega}(0)\tanh\tau\text{sech}\tau,$
$\hat{X}_{2\omega}(\tau)$ $=$ $\frac{1}{\sqrt{2}}(\tanh\tau+\tau\text{sech}^{2}\tau)\hat{X}_{\omega}(0)+\hat{X}_{2\omega}(0)\text{sech}^{2}\tau,$
$\hat{P}_{\omega}(\tau)$ $=$ $\hat{P}_{\omega}(0)\text{sech}\tau-\frac{1}{\sqrt{2}}(\sinh\tau+\tau\text{sech}\tau)\hat{P}_{2\omega}(0),$
$\hat{P}_{2\omega}(\tau)$ $=$ $\sqrt{2}\hat{P}_{\omega}(0)\tanh\tau+(1-\tau\tanh\tau)\hat{P}_{2\omega}(0).$ (71)
@@
$\langle\hat{X}^{2}_{j\omega}(\tau)\rangle=\langle\hat{X}^{2}_{j\omega}(0)\rangle S_{j\omega X},~~\text{and}~~\langle\hat{P}^{2}_{j\omega}(\tau)\rangle=\langle\hat{P}^{2}_{j\omega}(0)\rangle S_{j\omega P},$ (72)
where Li and Kumar (1994); Ou (1994):
$S_{\omega X}$ $=$ $(1-\tau\tanh\tau)^{2}\text{sech}^{2}\tau+2\tanh^{2}\tau\text{sech}^{2}\tau,$
$S_{2\omega X}$ $=$ $\frac{1}{2}(\tanh\tau+\tau\text{sech}^{2}\tau)^{2}+\text{sech}^{4}\tau,$
$S_{\omega P}$ $=$ $\frac{1}{2}(\sinh\tau+\tau\text{sech}\tau)^{2}+\text{sech}^{2}\tau,$
$S_{2\omega P}$ $=$ $2\tanh^{2}\tau+(1-\tau\tanh\tau)^{2}.$ (73)
For large $\tau$, note that, $S_{\omega X}\rightarrow 4\tau^{2}e^{-2\tau}$, $S_{2\omega X}\rightarrow 1/2$, $S_{\omega P}\rightarrow e^{2\tau}/8$, and $S_{2\omega P}\rightarrow\tau^{2}$ Ou (1994); Li and Kumar (1994); Suhara et al. (1996), suggesting the exponential enhancement in the noise of the phase quadrature. While the predictions are limited by the validity of the linearized equations that, for instance, assume a highly intense coherent field on the fundamental mode to start with, we can take these results to indicate that substantial squeezing can, in principle, be generated via the second harmonic generation in the fundamental mode. However, for the scenarios we are interested in, we have also excluded the decay, or damping, of the quasinormal modes, so our predictions are relevant for weakly damped quasinormal modes that are long-lived, applicable at timescales much smaller than the lifetime of quasinormal modes. Hence, the relevant degree of squeezing can be estimated based on the limit of small $\tau$, captured already by the perturbative expansion, Eq. (28) of the main text.